# Greedy Heuristics for Sampling-based Motion Planning in High-Dimensional State Spaces

### v1

@@ 3 Greedy Informed Set
Furthermore, it is bounded from above by the informed set, i.e., $\mathcal{X}_{\hat{f}_{\text{greedy}}}\supseteq\mathcal{X}_{\hat{f}}$ and will lead to smaller ellipsoidal regions compared to the original informed set, particularly in high-dimensional planning scenarios with complex solution paths. Consequently, sampling from this smaller subset will increase the likelihood of finding states that could belong to the omniscient set, thus facilitating faster convergence rates in improving the current solution path. However, because the greedy informed set is constructed based on the exploited knowledge of the existing solution path, similar to path biasing methods, it may or may not contain all the samples from the omniscient set, attempting to seek locally optimal solutions in certain planning contexts. Therefore, exploration into other homotopy classes, i.e., sampling from the informed set, remains essential for planners utilizing the greedy informed set to ensure global optimality and hold asymptotic optimality guarantees (see Section 4).
4 Greedy RRT* (G-RRT*)
@@
Greedy RRT* is an almost-surely asymptotically optimal path planning algorithm that build on top of RRT*. Since it is a RRT*-like algorithm, it rapidly explores the problem domain by building a tree through obstacle-free space with the help of random sampling. It also does the incremental rewiring of the growing tree to preserve the asymptotic optimality guarantees. However, instead of randomly sampling all over the problem domain throughout the entire planning process, G-RRT* uses the direct informed sampling technique (Gammell et al., 2014, 2018) to focus the search only on the promising regions of the problem state space once the initial solution is found. Specifically, it uses a greedy version of informed set introduced in Section 3 to focus the search with the exploited knowledge of the existing solution path.
@@
Finding the fast initial solution is important for G-RRT* to make use of all the benefits of informed sets since RRT* and its informed versions behave the same before finding the initial solution. One way to accelerate the finding of initial solution is to leverage on bi-directional search, for example, bi-directional version of RRT (Kuffner and LaValle, 2000) and its optimal variants (Klemm et al., 2015; Mashayekhi et al., 2020). Therefore, G-RRT* improves the initial path finding time by leveraging on the concepts of bi-directional search. It maintains two rapidly growing trees, one rooted in the start and one in the goal, to explore the problem state space and uses a greedy connect heuristic to guide the tree towards each other similar to RRT-Connect (Kuffner and LaValle, 2000).
@@ 3 $c_{\text{i}}\leftarrow\infty$;
7if $|X_{\text{soln}}|>0$ then
@@
9             $c_{\text{max}}\leftarrow-\infty$;
10             $\mathbf{x}_{\text{best}}\leftarrow\mathbf{x}_{\text{i}}$;
11             while $\text{Parent}(\mathbf{x}_{\text{best}})\neq\text{null}$ do
12                   $c_{\text{best}}\leftarrow\hat{f}(\mathbf{x}_{\text{best}})$;
13                   if $c_{\text{best}}>c_{\text{max}}$ then
14                         $c_{\text{max}}\leftarrow c_{\text{best}}$;
15
16                  $\mathbf{x}_{\text{best}}\leftarrow\text{Parent}(\mathbf{x}_{\text{best}})$;
17
@@ 7       $X_{\text{near}}\leftarrow\text{Near}\,(G=(V,E),\mathbf{x}_{\text{new}},r_{\text{rewire}})$;
8       $X_{\text{ancestors}}\leftarrow\text{Ancestors}\,(G=(V,E),X_{\text{near}},r_{\text{depth}})$;
@@ 9       $\mathbf{x}_{\text{min}}\leftarrow\mathbf{x}_{\text{nearest}}$;
10       foreach $\mathbf{x}_{\text{near}}\in(X_{\text{near}}\cup X_{\text{ancestors}})$ do
@@ 16      $E\leftarrow E\cup\{\mathbf{x}_{\text{min}},\mathbf{x}_{\text{new}}\}$;
17       foreach $\mathbf{x}_{\text{near}}\in(X_{\text{near}}\cup X_{\text{ancestors}})$ do
@@ 22                         $E\leftarrow E\cup\{(\mathbf{x}_{\text{new}},\mathbf{x}_{\text{near}})\}$;
29
30return $\textup{TRAPPED},\mathbf{x}_{\text{new}}$;
@@ 4.1 Notation
The state space of the path planning problem is denoted by $\mathcal{X}\in\mathbb{R}^{n}$ where $n\in\mathbb{N}$. Those states that are in collision with obstacles are denoted by $\mathcal{X}_{obs}\subseteq\mathcal{X}$ whereas valid collision-free states are denoted by $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$. A notation $\mathbf{x}\in\mathcal{X}$ represents a single state in the state space. The initial and goal states of the planning problem are denoted by $\mathbf{x}_{I}\in\mathcal{X}_{\text{free}}$ and $\mathbf{x}_{G}\in\mathcal{X}_{\text{free}}$ respectively.
@@
A tree $G=(V,E)$ is defined by a set of vertices $V$ and edges $E$. The start tree is denoted by $G_{a}=(V_{a},E_{a})$ and the goal tree is denoted by $G_{b}=(V_{b},E_{b})$. All the vertices in the start and goal trees $V_{a}$ and $V_{b}$ are directly associated with the valid regions of the state space $\mathcal{X}_{\text{free}}$. Edges in both trees $E\subset V\times V$ are directed with a source vertex $\mathbf{v}$ and a target vertex $\mathbf{w}$, represented by $(\mathbf{v},\mathbf{w})$.
@@
The cost of the connection from the source vertex $\mathbf{v}$ to the target vertex $\mathbf{w}$ in the graph $G$ is denoted by the scalar $c\in\mathbb{R}_{\geq 0}$ and its admissible heuristics is represented by $\hat{c}\in\mathbb{R}_{\geq 0}$ where $\forall{v,w}\in G,\,\hat{c}(\mathbf{v},\mathbf{w})\leq c(\mathbf{v},\mathbf{w})$. The true cost of the optimal path from the initial vertex to the specific vertex $\mathbf{x}$, also called, the cost-to-come of a vertex, is denoted by the function $g(\mathbf{x})=c(\mathbf{x}_{I},\mathbf{x})$ and its admissible estimates by $\hat{g}(\mathbf{x})=\hat{c}(\mathbf{x}_{I},\mathbf{x})$. The true cost of the optimal path from the vertex to the goal state, also called, the cost-to-go of a vertex is denoted by the function $h(\mathbf{x})=c(\mathbf{x}_{I},\mathbf{x})$ and its admissible estimates by $\hat{h}(\mathbf{x})=\hat{c}(\mathbf{x}_{I},\mathbf{x})$.
@@
The summation of these two heuristic estimates $\hat{g}(\mathbf{x})$ and $\hat{h}(\mathbf{x})$ defines the heuristic cost of the path from the start vertex $\mathbf{x}_{I}$ passing through a specific vertex $\mathbf{x}$ towards the goal vertex $\mathbf{x}_{G}$, i.e., $\hat{f}(\mathbf{x})=\hat{g}(\mathbf{x})+\hat{h}(\mathbf{x})$. This function $\hat{f}(\mathbf{x})$ also defines the informed set, i.e., the set of states whose heuristic estimates can provide better solutions than the current solution cost (Definition 6). The vertex with the maximum admissible heuristic cost in the path connected from the start vertex to the goal vertex is denoted by $\mathbf{x}_{max}$. This function $\hat{f}(\mathbf{x}_{max})$ defines the greedy informed set, i.e., the set of states whose heuristic estimates can provide better solutions than the heuristic cost of this greedily chosen state $\mathbf{x}_{max}$ in the current solution (Definition 7).
4.2 Greedy Informed Sampling

### v2

@@ III Greedy Informed Set
Furthermore, it is bounded from above by the informed set, i.e., $\mathcal{X}_{\hat{f}_{\text{greedy}}}\supseteq\mathcal{X}_{\hat{f}}$ and will lead to smaller ellipsoidal regions compared to the original informed set, particularly in high-dimensional planning scenarios with complex solution paths. Consequently, sampling from this smaller subset will increase the likelihood of finding states that could belong to the omniscient set, thus facilitating faster convergence rates in improving the current solution path. However, because the greedy informed set is constructed based on the exploited knowledge of the existing solution path, similar to path biasing methods, it may or may not contain all the samples from the omniscient set, potentially seeking locally optimal solutions in certain planning contexts. Therefore, exploration into other homotopy classes—i.e., sampling from the informed set—remains essential for planners utilizing the greedy informed set to ensure global optimality and hold asymptotic optimality guarantees (see Section IV).
Sampling informed sets is a necessary condition for ensuring asymptotic optimality in planning problems, as states from the optimal path always lie within these sets [31]. The greedy informed set also includes states from the optimal path, but only under the condition that the current path and the true optimal path lie within the same homotopy class (Theorem 1). This condition offers a reasonable approximation that relaxes some of the optimality constraints of informed sets while still providing faster convergence and optimality guarantees under certain conditions. The motivation behind the proposed greedy heuristic approximation stems from the observation that, for many holonomic planning problems seeking to minimize path length, this relaxation is often sufficient to produce high-quality solutions.
In the following, we will draw on concepts from homotopy theory [40] to formalize the optimality proof of the greedy informed set (Theorem 1) and present a counterexample illustrating cases where it fails to find the optimal solution (Theorem 2). First, we present the formal definition of the path homotopy (Definition 8).
Definition 8 (Homotopy class).
Two paths $\pi$ and $\pi^{*}$ are said to be path homotopic if they have the same initial and goal states $\mathbf{x}_{I}$ and $\mathbf{x}_{G}$ and if there exists a continuous function $f:[0,1]\times[0,1]\to\mathcal{X}$ (called a homotopy) such that:
  • $f(s,0)=\pi(s)$,
  • $f(s,1)=\pi^{*}(s)$,
  • $f(0,t)=\mathbf{x}_{I}$,
  • $f(1,t)=\mathbf{x}_{G}$,
for each $s\in[0,1]$ and each $t\in[0,1]$. Then the paths $\pi$ and $\pi^{*}$ are said to be in the same homotopy class.
The following lemma establishes a connection between the notions of homotopic paths and collision-free space.
Lemma 1 (Homotopic paths can be continuously deformed into each other).
Let $\pi$ and $\pi^{*}$ be two paths from the initial state, $\mathbf{x}_{I}$ to the goal state, $\mathbf{x}_{G}$, in the same homotopy class. Then, there exists a continuous deformation $f:[0,1]\times[0,1]\to\mathcal{X}_{\text{free}}$ such that:
$f(0,\cdot)=\pi\quad\text{and}\quad f(1,\cdot)=\pi^{*},$
and for each $t\in[0,1]$, $f(t,\cdot)$ is a valid collision-free path.
Proof.
By the definition of homotopy (Definition 8), two paths $\pi$ and $\pi^{*}$ in the same homotopy class can be continuously deformed into each other while having the same initial and goal states fixed. This deformation guarantees that for any $t\in[0,1]$, the path $f(t,\cdot)$ is feasible and stay within the collision-free configuration space $\mathcal{X}_{\text{free}}$. Hence, $\pi$ and $\pi^{*}$ are connected via a continuous deformation in $\mathcal{X}_{\text{free}}$. ∎
Next, we want to establish that this property still holds for subpaths of homotopic paths.
Lemma 2 (Subpaths of homotopic paths can also be continuously deformed into each other).
Let $\pi$ and $\pi^{*}$ be two paths from the initial state $\mathbf{x}_{I}$ to the goal state $\mathbf{x}_{G}$ that are homotopic in the same homotopy class. If $\pi_{[a,b]}$ is any subpath of $\pi$ and $\pi^{*}_{[c,d]}$ is any subpath of $\pi^{*}$, then there exists a continuous deformation that allows $\pi_{[a,b]}$ to be deformed into $\pi^{*}_{[c,d]}$.
Proof.
Consider the intervals $[a,b]$ and $[c,d]$ corresponding to the subpaths $\pi_{[a,b]}$ and $\pi^{*}_{[c,d]}$, respectively. The function $g:[0,1]\times[a,b]\to\mathcal{X}_{\text{free}}$ can be defined as follows:
$g(t,s)=f(t,s),\,s\in[a,b]$
At $t=0$, the deformation begins with:
$g(0,s)=f(0,s)=\pi(s)\quad\text{for }s\in[a,b],$
and at $t=1$, the deformation concludes with:
$g(1,s)=f(1,s)=\pi^{*}(s)\quad\text{for }s\in[c,d].$
Thus, the function $g(t,s)$ represents a continuous transformation that allows $\pi_{[a,b]}$ to be deformed into $\pi^{*}_{[c,d]}$. Since $f(t,\cdot)$ is a valid collision-free path for every $t$ (Lemma 1), the deformation $g(t,s)$ remains within $\mathcal{X}_{\text{free}}$ throughout the deformation process. Hence, the conclusion is reached that any subpath of $\pi$ can be continuously deformed into any subpath of $\pi^{*}$ as long as both paths are homotopic. ∎
Next, we define the geodesic path on an ellipsoid (Definition 9), as it is crucial for establishing the optimality of the greedy informed set, and it is illustrated in Figure 4.
Figure 4: Illustration of the geodesic path between two states, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, on an ellipsoid, $\mathcal{E}$, in $\mathbb{R}^{2}$. The geodesic path, $\pi^{*}(\mathcal{E})$, is shown with a dashed line, while a higher-cost path, $\pi$, which passes outside the ellipsoid, is depicted with a solid line. The ellipsoid $\mathcal{E}$ is partially highlighted in the gray-shaded region.
Definition 9 (Geodesic path).
Let $\mathcal{E}$ be an ellipsoid in $\mathbb{R}^{n}$. Let $\partial\mathcal{E}$ be the boundary of $\mathcal{E}$, and let $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ be the two states on $\partial\mathcal{E}$. Define $\pi^{*}$ as the path along the boundary of $\mathcal{E}$ from $\mathbf{x}_{1}$ to $\mathbf{x}_{2}$, and let $\pi$ be the path that traverses outside $\mathcal{E}$ between these two points. We define the geodesic path of $\mathcal{E}$ between $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ as the direct path along the boundary of $\mathcal{E}$ that has minimal cost compared to all other paths that pass outside the ellipsoid between the same two endpoints (Figure 4), i.e.,
$\pi^{*}(\mathcal{E})=\underset{\pi\in\Sigma}{\operatorname{\arg\min}}\Big{\{}c(\pi)\mid\pi(0)=\mathbf{x}_{1},\ \pi(1)=\mathbf{x}_{2},\Big{.}$
.$\Big{.}\forall s\in[0,1],\ \pi(s)\in\mathcal{X}_{\text{free}}\cap\partial\mathcal{E}\setminus\mathcal{E}\Big{\}}.$
We now present the proof that the optimal path lies within the greedy informed set, given that both the optimal path and the current solution path are in the same homotopy class (Theorem 1). First, it demonstrates a case where the optimal solution lies outside the greedy hyperellipsoid and shows that this violates the optimality condition.
Figure 5: Illustration of the current solution path $\pi$ and another path $\pi^{*}$ in the same homotopy class, both with the same initial and goal states, $\mathbf{x}_{I}$ and $\mathbf{x}_{G}$. The greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}$ (depicted as a dashed gray ellipse) is constructed based on the maximum heuristic cost along $\pi$. The path $\pi^{*}$ intersects the boundary of the greedy informed set at states $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, passing through a state $\mathbf{x}_{3}$, which lies outside $\mathcal{X}_{\hat{f}_{\text{greedy}}}$.
Theorem 1 (Inclusion of the optimal path in the greedy informed set within the same homotopy class).
For a given path $\pi$, the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi)$ contains the optimal path $\pi^{*}$ if $\pi^{*}$ is in the same homotopy class as $\pi$.
Proof.
Proof of Theorem 1 follows directly from Lemma 2. We begin by constructing an example planning scenario, as depicted in Figure 5, with the details provided below.
Let $\pi$ be the current solution path in $\mathcal{X}_{\text{free}}$ from the initial state $\mathbf{x}_{I}$ to the goal state $\mathbf{x}_{G}$. Denote $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ as two states along the path $\pi$ that have the same maximum heuristic cost, i.e., $\hat{f}(\mathbf{x}_{1})=\hat{f}(\mathbf{x}_{2})=\hat{f}(\mathbf{x}_{\text{max}})$. The path $\pi$ is constructed such that any state $\mathbf{x}$ lying between $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ along the segment of $\pi$ has the same heuristic cost, i.e.,
$\forall\mathbf{x}\in\mathcal{Z},\,\hat{f}(\mathbf{x})=\hat{f}(\mathbf{x}_{\text{max}})$
where $\mathcal{Z}:=\{\mathbf{x}\in\pi^{\prime}\mid\pi^{\prime}(0)=\mathbf{x}_{1},\,\pi^{\prime}(1)=\mathbf{x}_{2},\,\pi^{\prime}\subseteq\pi\}$. By Definition 7, the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi)$ is constructed based on $\mathbf{x}_{\text{max}}$ of $\pi$. We aim to show that the optimal path $\pi^{*}$ also lies inside $\mathcal{X}_{\hat{f}_{\text{greedy}}}$ constructed by $\pi$ if both $\pi$ and $\pi^{*}$ are in the same homotopy class.
Theorem 1 is now proven by contradiction. Suppose there exists a state $\mathbf{x}$, denoted as $\mathbf{x}^{*}$, along $\pi^{*}$ (where $\pi^{*}$ is the optimal path) that is not included in the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi)$, i.e.,
$\forall\mathbf{x}\in\mathcal{X}_{\text{free}},\,\exists\mathbf{x}\in\pi^{*}\quad\text{s.t.}\quad\mathbf{x}\notin\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi).$ (12)
In this case, $\pi^{*}$ must leave the ellipsoid at some point and re-enter it, since both the initial state $\mathbf{x}_{I}$ and the goal state $\mathbf{x}_{G}$ are inside $\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi)$. Consequently, there will be at least two intersection points between $\pi^{*}$ and the boundary of $\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi)$. For simplicity, let the states $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ be the two intersection points, and $\mathbf{x}_{3}$ represent the state outside the hyperellipsoid.
Since both $\pi$ and $\pi^{*}$ are homotopic paths, the subpath $\pi^{*}_{\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3}}$ of $\pi^{*}$ can be continuously deformed onto the subpath $\pi_{\mathbf{x}_{1}\mathbf{x}_{2}}$ of $\pi$ without colliding with any obstacles (Lemma 2). By Definition 9, the cost of the deformed geodesic subpath $\pi^{*}_{\mathbf{x}_{1}\mathbf{x}_{2}}$ is lower than the cost of the subpath $\pi^{*}_{\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3}}$:
$c(\pi^{*}_{\mathbf{x}_{1}\mathbf{x}_{2}})<c(\pi^{*}_{\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3}}).$ (13)
Therefore, connecting directly to $\mathbf{x}_{2}$ is shorter than passing through $\mathbf{x}_{3}$. According to the definition of an optimal path, we know that any subpath of it must also be optimal [41]. However, by (13), the cost of $\pi^{*}$ is higher than the cost of the deformed path $\pi_{\mathbf{x}_{I}\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{G}}$. This implies that traversing through $\mathbf{x}_{3}$ incurs a higher cost compared to connecting directly from $\mathbf{x}_{1}$ to $\mathbf{x}_{2}$ along the ellipsoid, which contradicts the optimality of $\pi^{*}$. Thus, the initial assumption that $\mathbf{x}^{*}$ is not in the greedy informed set by (12) must be false and proves Theorem 1. Therefore, the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}(\pi)$ must include all the states along the optimal path $\pi^{*}$ when $\pi$ and $\pi^{*}$ are in the same homotopy class. ∎
While Theorem 1 establishes that the optimal path is guaranteed to be included in the greedy informed set when it shares the same homotopy class as the current solution path, this does not hold when the paths are homotopically distinct. In fact, there are cases where the greedy informed set may fail to include the optimal path precisely because of differences in homotopy classes. This leads us to explore scenarios where the optimal path is excluded from the greedy informed set (Theorem 2).
Theorem 2 (Scenarios where the greedy informed set does not contain the optimal path).
There exist scenarios in which the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}$ does not contain the optimal path $\pi^{*}$ due to their being in different homotopy classes.
Proof.
This proof is based on a counterexample and follows directly from Theorem 1. Consider a planning problem scenario in a maze-like environment with many homotopy classes between the initial state $\mathbf{x}_{I}$ and the goal state $\mathbf{x}_{G}$, as illustrated in Figure 6. Let $\pi_{i}$ be a feasible path that follows a narrow corridor, navigating through obstacles by making several sharp turns. Let $\pi^{*}$ be the optimal path in a different homotopy class, which avoids the narrow corridor. Instead, it circumvents the obstacles entirely, avoiding sharp turns. By construction, the greedy informed set is constrained by the geometry of the environment, specifically by the obstacles that force $\pi_{i}$ through the narrow passage.
Since $\pi_{i}$ and $\pi^{*}$ belong to different homotopy classes, the path $\pi_{i}$ cannot be continuously deformed into $\pi^{*}$ without intersecting the obstacles (Lemma 1). Theorem 1 guarantees that the optimal path $\pi^{*}$ will only be included in the greedy informed set if both $\pi_{i}$ and $\pi^{*}$ are homotopic. However, since $\mathcal{X}_{\hat{f}_{\text{greedy}}}$ is localized to states near the narrow passage, it excludes the optimal path $\pi^{*}$, which circumvents the obstacles. Therefore, in scenarios where the current feasible solution path and the optimal path belong to different homotopy classes, the greedy informed set may fail to capture the optimal path. ∎
Figure 6: Illustration of the sub-optimality of greedy informed set in an example planning scenario. The $L^{2}$ greedy informed set, $\mathcal{X}_{\hat{f}_{\text{greedy}}}$, highlighted in a dashed ellipse, is constructed based on the current tortuous solution path, $\pi_{i}$. The true optimal path, $\pi^{*}$, is depicted as a dotted line. Here, the greedy informed set does not contain all the vertices that can improve the current solution cost (i.e., those that contribute to approaching $\pi^{*}$ and lie outside of the hyperellipsoid) due to the nature of its greedy exploitation.
This highlights the limitations of the greedy informed set in certain planning problems. When the optimal path belongs to a different homotopy class than the current solution path, the greedy informed set cannot guarantee the inclusion of the optimal path, which can lead to suboptimal performance. Therefore, understanding this behavior is crucial, as it allows for the appropriate adjustment of the sampling rate from this set based on the distribution of homotopy classes and the type of planning problem being addressed. One effective approach to mitigate this issue is to also sample from the informed set with a certain probability (Section IV), which helps control the balance between exploration and exploitation within the planning algorithm.
IV Greedy RRT* (G-RRT*)
@@
Greedy RRT* is an almost surely asymptotically optimal path planning algorithm that builds on top of RRT*. Since it is an RRT*-like algorithm, it rapidly explores the problem domain by building a tree through obstacle-free space using random sampling. It also does the incremental rewiring of the growing tree to preserve the asymptotic optimality guarantees. However, rather than sampling randomly throughout the problem domain for the entire planning process, G-RRT* uses the direct informed sampling technique [12, 31] to focus the search on the promising regions of the state space once the initial solution is found. Specifically, it uses a greedy version of the informed set, introduced in Section III, to focus the search by exploiting knowledge of the existing solution path.
@@
Finding a fast initial solution is important for G-RRT* to make use of all the benefits of informed sets, as RRT* and its informed variants behave the same until the initial solution is found. One way to accelerate this process is by utilizing bi-directional search, such as the bi-directional version of RRT [13] and its optimal variants [21, 42]. Therefore, G-RRT* improves the initial path finding time by leveraging the concepts of bi-directional search. It maintains two rapidly growing trees—one rooted at the start and one at the goal—to explore the problem’s state space and uses a greedy connect heuristic to guide the trees toward each other, similar to RRT-Connect [13] algorithm. The complete algorithm is presented in Algorithms 15, with modifications to the bi-directional versions of RRT and RRT* variants highlighted in red.
@@ 3 $c_{\text{i}}\leftarrow\infty$;
7      $c_{\text{i}}\leftarrow\underset{\mathbf{x}\in X_{\text{soln}}}{\operatorname{\min}}\left\{g(\mathbf{x})\right\}$;
8
@@
10
11            if $c_{\text{i}}<c_{\text{min}}$ then
12                   $c_{\text{min}}\leftarrow c_{\text{i}}$;
13                   $c_{\text{max}}\leftarrow\underset{\mathbf{x}\in X_{\text{soln}}}{\operatorname{\max}}\left\{\hat{f}(\mathbf{x})\right\}$;
14
@@ 8       $\mathbf{x}_{\text{min}}\leftarrow\mathbf{x}_{\text{nearest}}$;
9       foreach $\mathbf{x}_{\text{near}}\in X_{\text{near}}$ do
@@ 15      $E\leftarrow E\cup\{\mathbf{x}_{\text{min}},\mathbf{x}_{\text{new}}\}$;
16       foreach $\mathbf{x}_{\text{near}}\in X_{\text{near}}$ do
@@ 21                         $E\leftarrow E\cup\{(\mathbf{x}_{\text{new}},\mathbf{x}_{\text{near}})\}$;
28
29return $\textup{TRAPPED},\mathbf{x}_{\text{new}}$;
@@ IV-A Notation
The state space of the path planning problem is denoted by $\mathcal{X}\in\mathbb{R}^{n}$, where $n\in\mathbb{N}$. States that are in collision with obstacles are represented by $\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}$, while valid, collision-free states are denoted by $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$. The notation $\mathbf{x}\in\mathcal{X}$ represents a single state in the state space. The initial and goal states of the planning problem are denoted by $\mathbf{x}_{I}\in\mathcal{X}_{\text{free}}$ and $\mathbf{x}_{G}\in\mathcal{X}_{\text{free}}$, respectively.
@@
A tree $G=(V,E)$ is defined by a set of vertices $V$ and edges $E$. The start tree is denoted by $G_{a}=(V_{a},E_{a})$, and the goal tree is denoted by $G_{b}=(V_{b},E_{b})$. All vertices in the start and goal trees, $V_{a}$ and $V_{b}$, are directly associated with the valid regions of the state space $\mathcal{X}_{\text{free}}$. Edges in both trees, $E\subset V\times V$, are directed and consist of a source vertex $\mathbf{v}$ and a target vertex $\mathbf{w}$, represented as $(\mathbf{v},\mathbf{w})$.
@@
The cost of the connection from the source vertex $\mathbf{v}$ to the target vertex $\mathbf{w}$ in the graph $G$ is denoted by the scalar $c\in\mathbb{R}_{\geq 0}$, and its admissible heuristic is represented by $\hat{c}\in\mathbb{R}_{\geq 0}$, where $\forall{v,w}\in G,\,\hat{c}(\mathbf{v},\mathbf{w})\leq c(\mathbf{v},\mathbf{w})$. The true cost of the optimal path from the initial vertex to a specific vertex $\mathbf{x}$, also called the cost-to-come of a vertex, is denoted by the function $g(\mathbf{x})=c(\mathbf{x}_{I},\mathbf{x})$, and its admissible estimate is $\hat{g}(\mathbf{x})=\hat{c}(\mathbf{x}_{I},\mathbf{x})$. The true cost of the optimal path from a vertex to the goal state, also called the cost-to-go of a vertex, is denoted by the function $h(\mathbf{x})=c(\mathbf{x},\mathbf{x}_{G})$, and its admissible estimate is $\hat{h}(\mathbf{x})=\hat{c}(\mathbf{x},\mathbf{x}_{G})$.
@@
The summation of the two heuristic estimates, $\hat{g}(\mathbf{x})$ and $\hat{h}(\mathbf{x})$, defines the heuristic cost of the path from the start vertex $\mathbf{x}_{I}$, passing through a specific vertex $\mathbf{x}$, towards the goal vertex $\mathbf{x}_{G}$, i.e., $\hat{f}(\mathbf{x})=\hat{g}(\mathbf{x})+\hat{h}(\mathbf{x})$. This function, $\hat{f}(\mathbf{x})$, also defines the informed set, which is the set of states whose heuristic estimates can provide better solutions than the current solution cost (Definition 6). The vertex with the maximum admissible heuristic cost along the path from the start vertex to the goal vertex is denoted by $\mathbf{x}_{\text{max}}$. The function $\hat{f}(\mathbf{x}_{\text{max}})$ defines the greedy informed set, which is the set of states whose heuristic estimates can provide better solutions than the heuristic cost of this greedily chosen state, $\mathbf{x}_{\text{max}}$, in the current solution (Definition 7).
IV-B Greedy Informed Sampling